# Scaling FP8 training to trillion-token LLMs

### v1

@@ Abstract
We train, for the first time, large language models using FP8 precision on datasets up to 2 trillion tokens — a 20-fold increase over previous limits. Through these extended training runs, we uncover critical instabilities in FP8 training that were not observable in earlier works with shorter durations. We trace these instabilities to outlier amplification by the SwiGLU activation function. Interestingly, we show, both analytically and empirically, that this amplification happens only over prolonged training periods, and link it to a SwiGLU weight alignment process. To address this newly identified issue, we introduce Smooth-SwiGLU, a novel modification that ensures stable FP8 training without altering function behavior. We also demonstrate, for the first time, FP8 quantization of both Adam optimizer moments. Combining these innovations, we successfully train a 7B parameter model using FP8 precision on 256 Intel Gaudi2 accelerators, achieving on-par results with the BF16 baseline while delivering up to a throughput improvement.
@@ 1 Introduction
In this paper, we present advancements in FP8 training for LLMs, successfully scaling to datasets of up to 2 trillion tokens - a 20-fold increase over previous limits. This leap in scale has revealed critical instabilities in FP8 training that were not observable in earlier, shorter-duration studies. Through rigorous analysis, we trace these instabilities to a previously unidentified phenomenon: the amplification of outliers by the SwiGLU activation function (Shazeer (2020a)), which becomes pronounced only after extended training periods. The severity of this issue is illustrated in Fig. 2(a), where we present the training loss of Llama2 7B using FP8 precision. The graph clearly shows a dramatic divergence caused by outliers after processing 220B tokens - twice the dataset size explored in previous work (Peng et al. (2023)).
@@ 2 Background and Challenges of FP8 Training in LLMs
Recent work by Peng et al. (2023) demonstrated the first empirical results of training LLMs using FP8 format up to 100 billion tokens, validating the potential of FP8 for large-scale training. Our research extends this work, successfully scaling FP8 training to datasets of up to 2 trillion tokens. We introduce novel techniques to address the challenges of FP8’s limited dynamic range and the instabilities that emerge in extended training scenarios. Our approach not only overcomes these limitations but also achieves substantial improvements in memory usage and training speed, demonstrating the viability of FP8 training for truly large-scale LLM development.
@@ 3 Outlier Amplification in Large-Scale FP8 Training
The presence of outliers has been observed in numerous studies (Yang et al., 2024), particularly in the activations during inference. These outliers can significantly impact the stability and performance of the model, as they introduce extreme values that are difficult to manage within the limited dynamic range of reduced-precision formats like FP8. Our work reveals that these outliers become particularly prominent in the later stages of training large language models (LLMs) with large-scale datasets.
As shown in Fig. 1, these outliers emerge only after processing approximately 200 billion tokens during training. This phenomenon poses significant challenges to maintaining numerical stability and model performance, especially when using delayed scaling methods that assume consistency across iterations. The sudden appearance of these outliers disrupts the statistical assumptions underlying FP8 training techniques, potentially leading to instability or divergence in the training process.
@@ 4.1 SwiGLU Structure
Let be the input vector from the previous layer. The SwiGLU neuron is defined as follows for two weight vectors :
@@
While other standard neuron types, such ReLU, GeLU, and Swish at most linear at large input magnitudes (i.e., ), the SwiGLU neuron is quadratic and can reach much larger values (and cause very strong outliers) if and are sufficiently aligned (e.g., if and then ). As we show next, precisely such alignment happens during training for neurons with sufficiently large inputs.
@@ 4.2 Theoretical analysis of weight correlation in SwiGLU
To show this, We assume the SwiGLU neuron is embedded in a neural network with parameters. The rest of the parameters in the network are denoted by . We train the neural network with some regularization,
@@
Using the chain rule we have:
Note where we defined and with a slight abuse of notation, we suppressed the dependence of and on . Given the assumption
@@
Now, defining , the above equations become:
@@
Thus, with regularization and sufficient training, the weight vectors and must converge to either identical or opposite directions, i.e., or — if holds for all samples . Since decays exponentially fast as increases, this simply means that neuron inputs are not too small. This can happen in the case in which is sufficiently large, and . The latter condition is the generic case when we fit a neural network to a large dataset of size , where (i.e. we are in an under-parameterized regime). . Interestingly, this alignment phenomenon happens due to regularization, even if it is very weak. In fact, weak regularization will lead to larger weight norms and will strengthen this effect.
@@ 4.3 Observing weight correlation growth and training instability
The process of weight alignment and its impact on FP8 training develops precisely as our theory predicts. As training progresses, input magnitudes in certain channels grow, eventually exceeding a critical threshold. When this occurs, weight vectors and in these channels begin to align rapidly. Figure 2b illustrates this evolution: initially, the correlation between and is low, but it increases drastically between 125B and 210B tokens. Simultaneously, we observe a significant increase in the norm of these weights. This combination of high correlation and increased norm creates ideal conditions for generating extreme activation values, or “spikes.”
@@ 4.4 Smooth-SwiGlu
As demonstrated earlier, the SwiGLU activation function can lead to outliers in the input of the last linear layer of the MLP component. These outliers pose a significant challenge when using FP8 precision with delayed scaling, which relies on the assumption that statistical properties remain consistent across layers. The sudden spike in value caused by SwiGLU disrupts this continuity, leading to instability in the training process. In Fig. 3 we demonstrate that disabling the quantization of the last linear layer in the MLP component (output of SwiGLU) allows Llama2 FP8 to successfully converge with large datasets, addressing the previously observed divergence issues.
@@ 5 FP8 Optimizer
In our experiments, presented in Fig. 5 we show that while the first moment is able to converge with E4M3, the second moment requires a wider dynamic range and is able to converge only with E5M2 format. In Table 1 we compare the proposed quantization scheme for the optimizer moments with the one presented in Peng et al. (2023). Our scheme shows, for the first time, the ability to quantize both moments with standard FP8 formats.
@@ 6.2 Results
In Fig. 6 we show the training loss of Llama2 over 300B tokens with the proposed scheme, which include the use of Smooth SwiGLU (Section 4.4) + FP8 quantization of both Adam moments (Section 5). Notice we are able to overcome the divergence point of standard FP8 training . Additional training steps will be added in the next version of the manuscript. The FP8 model was trained using the standard format (Micikevicius et al., 2022) which include E4M3 for the forward phase and E5M2 for the backward phase.
@@
Table 2 compares the zero-shot performance (accuracy and perplexity) on downstream tasks between the BF16 baseline and our FP8 model with Smooth-SwiGLU and FP8 optimizer. The results demonstrate that our FP8 approach achieves on-par performance with the BF16 baseline across all tested metrics.
@@
Configuration | Micro BS | Status | Throughput |
(Samples/sec) | |||
BF16 | 1 | 12.65 | |
FP8 + SwiGLU output in BF16 | 1 | Converge | 16.07 (+ 27.04 %) |
FP8 + Smooth SwiGLU | 1 | Converge | 16.89 (+33.52 %) |
FP8 | 1 | Diverge | 17.34 (+37.08 %) |
@@
In Table 4 we present the memory reduction achieved by changing the optimizer moments from standard FP32 to FP8. As can be seen, we can reduce the memory consumption by ,
@@ 7 Conclusions
Our proposed method, combining Smooth-SwiGLU and FP8 optimizer moments, achieved comparable performance to BF16 baselines on downstream tasks while providing significant throughput improvementsd. This approach successfully overcome the divergence challenges typically encountered in standard FP8 training on large datasets.

### v2

@@ Abstract
We train, for the first time, large language models using FP8 precision on datasets up to 2 trillion tokens — a 20-fold increase over previous limits. Through these extended training runs, we uncover critical instabilities in FP8 training that were not observable in earlier works with shorter durations. We trace these instabilities to outlier amplification by the SwiGLU activation function. Interestingly, we show, both analytically and empirically, that this amplification happens only over prolonged training periods, and link it to a SwiGLU weight alignment process. To address this newly identified issue, we introduce Smooth-SwiGLU, a novel modification that ensures stable FP8 training without altering function behavior. We also demonstrate, for the first time, FP8 quantization of both Adam optimizer moments. Combining these innovations, we successfully train a 7B parameter model using FP8 precision on 256 Intel Gaudi2 accelerators, achieving on-par results with the BF16 baseline while delivering up to a throughput improvement. A reference implementation is supplied in https://github.com/Anonymous1252022/Megatron-DeepSpeed.
@@ 1 Introduction
In this paper, we present advancements in FP8 training for LLMs, successfully scaling to datasets of up to 2 trillion tokens — a 20-fold increase over previous limits. This leap in scale has revealed critical instabilities in FP8 training that were not observable in earlier, shorter-duration studies. Through rigorous analysis, we trace these instabilities to a previously unidentified phenomenon: the amplification of outliers by the SwiGLU activation function (Shazeer (2020a)), which becomes pronounced only after extended training periods. The severity of this issue is illustrated in Fig. 2(a), where we present the training loss of Llama2 7B using FP8 precision. The graph clearly shows a dramatic divergence caused by outliers after processing 220B tokens - twice the dataset size explored in previous work (Peng et al. (2023)).
@@ 2 Background and Challenges of FP8 Training in LLMs
Recent work by Peng et al. (2023) demonstrated the first empirical results of training LLMs using FP8 format up to 100 billion tokens, validating the potential of FP8 for large-scale training, but less applicable in real-world scenarios that usually require larger training. Our research extends this work, successfully scaling FP8 training to datasets of up to 2 trillion tokens. We introduce novel techniques to address the challenges of FP8’s limited dynamic range and the instabilities that emerge in extended training scenarios. Our approach not only overcomes these limitations but also achieves substantial improvements in memory usage and training speed, demonstrating the viability of FP8 training for truly large-scale LLM development.
@@ 3 Outlier Amplification in Large-Scale FP8 Training
The presence of outliers has been observed in numerous studies (Yang et al., 2024; Bondarenko et al., 2023), particularly in the activations during inference. One of the previously solution presented to confront with these outliers for inference, is to apply rotation to the activations (Liu et al., 2024). These outliers can significantly impact the stability and performance of the model, as they introduce extreme values that are difficult to manage within the limited dynamic range of reduced-precision formats like FP8. Our work reveals that these outliers become particularly prominent in the later stages of training large language models (LLMs) with large-scale datasets.
As shown in Fig. 1, these outliers emerge only after processing approximately 200 billion tokens during training. This phenomenon poses significant challenges to maintaining numerical stability and model performance, especially when using methods that assume consistency across iterations. The sudden appearance of these outliers, which are crucial for model performance (Sun et al., 2024), disrupts the statistical assumptions underlying FP8 training techniques, potentially leading to instability or divergence in the training process.
@@ 4.1 SwiGLU Structure
Let be the input vector from the previous layer. For two weight vectors , the SwiGLU neuron is defined as follows
@@
While other standard neuron types, such ReLU, GeLU, and Swish at most linear at large input magnitudes (i.e., ), the SwiGLU activation is quadratic and can reach much larger values (and cause very strong outliers) if and are sufficiently aligned (e.g., if and then ). As we show next, precisely such alignment happens during training for neurons with sufficiently large inputs.
@@ 4.2 Theoretical analysis of weight correlation in SwiGLU
To show this, we assume the SwiGLU neuron is embedded in a neural network with parameters. The rest of the parameters in the network are denoted by . We train the neural network with some regularization
@@
Using the chain rule we obtain the following two equations
where we defined and with a slight abuse of notation, we suppressed the dependence of and on . Given the assumption
@@
Now, defining , the above equations become
@@
Note this result holds also if we replace the swish activation in SwiGLU with other GLU variants (Shazeer, 2020a), since we did not use any specific properties of the Swish. Thus, practically, with regularization and sufficient training, the weight vectors and must converge to either identical or opposite directions, i.e., or — if is typically small. Since decays exponentially fast as increases, this simply means that the neuron inputs are typically not too small. This can happen in the case in which is sufficiently large, and typically . One should expect the latter condition to be the generic case when we fit a neural network to a large dataset of size , where (i.e., we are in an under-parameterized regime) and zero loss is not reachable—since then the neural network does not have spare capacity to set specific neuron inputs to zero (in addition to fitting the data). And indeed, we observe (Fig. 9 in the Appendix) that after training for of the tokens, in the outlier neuron. Interestingly, this alignment phenomenon occurs due to regularization, even if it is very weak. In fact, weak regularization will lead to larger weight norms, strengthening this effect.
@@ 4.3 Observing weight correlation growth and training instability
In Fig. 2 we see the process of weight alignment and its impact on FP8 training develops precisely as our theory predicts. As training progresses, in the outlier channel grows, eventually exceeding a critical threshold satisfying our Theorem’s assumption. Thus, the weight vectors and in these channels begin to align rapidly—i.e. the correlation between and is initially low, but then it increases drastically between 125B and 210B tokens. Interestingly, it seems this alignment happens simultaneously with further norm growth. This combination of high correlation and increased weight norm creates ideal conditions for generating extreme activation values, or “spikes.”
@@ 4.4 Smooth-SwiGlu
As demonstrated earlier, the SwiGLU activation function can lead to outliers in the input of the last linear layer of the MLP component. These outliers pose a significant challenge when using FP8 precision with delayed scaling, which relies on the assumption that statistical properties remain consistent across layers. The sudden spike in value caused by SwiGLU disrupts this continuity, leading to instability in the training process. In Fig. 3 we demonstrate that disabling the quantization of the last linear layer in the MLP component (output of SwiGLU) allows Llama2 FP8 to successfully converge with large datasets, addressing the previously observed divergence issues. This shows that other components in Llama architecture, such as RMS Norm or MHA are not the cause of instability in FP8 training.
@@ 5.2 Methodology
In our experiments, presented in Fig. 5 we show that while the first moment is able to converge with E4M3, the second moment, which estimates the square of the gradients, requires a wider dynamic range and is able to converge only with E5M2 format. In Table 1 we compare the proposed quantization scheme for the optimizer moments with the one presented in Peng et al. (2023). Our scheme shows, for the first time, the ability to quantize both moments with standard FP8 formats.
@@ 6.2 Results
In Fig. 6 we show the training loss of Llama2 with the proposed scheme, which includes the use of Smooth SwiGLU (Section 4.4) + FP8 quantization of both Adam moments (Section 5). Notice we can overcome the divergence point of standard FP8 training. The FP8 model was trained using the standard format (Micikevicius et al., 2022) which includes saving a high precision weight matrix and quantization to E4M3 for the forward phase and E5M2 for the backward phase with delayed scaling, similar to Nvidia’s transformer Engine. The model was trained on 256 Intel Gaudi2 over 15 days.
@@
Table 2 compares the zero-shot performance (accuracy and perplexity) on downstream tasks between the BF16 baseline and our FP8 model. The results demonstrate that our FP8 approach achieves on-par performance with the BF16 baseline across all tested metrics.
@@
Configuration | Micro BS | Status | Throughput | TFLOPS |
(Samples/sec) | ||||
BF16 | 1 | 12.65 | 311 | |
FP8 + SwiGLU output in BF16 | 1 | Converge | 16.07 (+ 27.04 %) | 396 |
FP8 + Smooth SwiGLU | 1 | Converge | 16.89 (+33.52 %) | 417 |
FP8 | 1 | Diverge | 17.34 (+37.08 %) | 428 |
@@
In Table 4 we present the memory reduction achieved by changing the optimizer moments from standard FP32 to FP8. Moreover, we reduce the master weight to FP16, as shown in Peng et al. (2023). As can be seen, we can reduce the memory consumption by ,
@@ 7 Conclusions
Our proposed method, combining Smooth-SwiGLU and FP8 optimizer moments, achieved comparable performance to BF16 baselines on downstream tasks while providing significant throughput improvements. This approach successfully overcome the divergence challenges typically encountered in standard FP8 training on large datasets.
Reproducibility
The abstract of the paper provides a link to an anonymous GitHub repository (https://github.com/Anonymous1252022/Megatron-DeepSpeed) containing all the code and necessary details for reproducing the experiments.
Ethics
LLMs require immense computational resources during training, which contributes significantly to carbon emissions. This environmental cost has become a growing concern in the field of AI. The use of low-precision formats like FP8 offers a promising solution, as it significantly reduces the computational overhead without sacrificing model accuracy. By adopting FP8 for training, not only can we enhance training efficiency, but we can also mitigate the carbon footprint associated with large-scale LLM training, paving the way for more sustainable AI development.
Acknowledgements
The research of DS was Funded by the European Union (ERC, A-B-C-Deep, 101039436). Views and opinions expressed are however those of the author only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency (ERCEA). Neither the European Union nor the granting authority can be held responsible for them. DS also acknowledges the support of the Schmidt Career Advancement Chair in AI.
@@ A.1 Training instability - additional data
A.2 Performance gain on Nvidia GPUs
Configuration | Micro BS | Status | Throughput | TFLOPS |
(Samples/sec) | ||||
BF16 | 1 | 3.22 | 76 | |
FP8 + SwiGLU output in BF16 | 1 | Converge | 4.11 (+ 27.6 %) | 96.9 |
FP8 + Smooth SwiGLU | 1 | Converge | 4.32 (+34.16 %) | 101.9 |
FP8 | 1 | Diverge | 4.43 (+37.58 %) | 104.5 |
A.3 Smooth-SwiGLU study
In Fig. 10 we show a study of the effect of Smooth-SwiGLU on BF16 training with different LR. Smooth-SwiGLU allows a smoother training curve even with BF16 training. Moreover, it enables training to lower loss values, especially when using a larger LR. (Fig. 11)
A.4 FP8 without SwiGLU activation function
In Fig. 12 we show FP8 training on c4 dataset of GPT3 model, which include GeLU activation function. Notice, in this scenario no training stability issues were observed.